# Intuitive Fine-Tuning: Towards Unifying SFT and RLHF into a Single Process

### v1

@@ Abstract
Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) are two fundamental processes for enhancing the capabilities of Language Models (LMs) post pre-training, aligning them better with human preferences. Although SFT advances in training efficiency, RLHF delivers better alignment, thus they are often combined. However, common practices simply apply them sequentially without unifying their optimization targets, resulting in trade-offs between fitting different objectives. This approach ignores the opportunities to bridge the paradigm gap and take the strengths from both. To obtain a unified understanding, we interpret SFT and RLHF with two sub-processes — Preference Estimation and Transition Optimization — defined at token level within the Markov Decision Process (MDP) framework. This modeling shows that SFT is only a specialized case of RLHF with inferior estimation and optimization. RLHF evaluates the quality of model’s entire generated answer, whereas SFT only scores predicted tokens based on preceding tokens from target answers. Therefore, SFT overestimates the ability of model, leading to inferior optimization. Building on this view, we introduce Intuitive Fine-tuning (IFT) to integrate SFT and RLHF into a single process. IFT captures LMs’ intuitive sense of the entire answers through a temporal residual connection, but it solely relies on a single policy and the same volume of non-preference-labeled data as SFT. Our experiments show that IFT performs comparably or even superiorly to sequential recipes of SFT and some typical alignment methods across several tasks, particularly those requires generation, reasoning, and fact-following abilities. An explainable Frozen Lake game further validates the effectiveness of IFT.
Intuitive Fine-Tuning:
Towards Unifying SFT and RLHF into a Single Process
Ermo Hua1, Biqing Qi1,2, Kaiyan Zhang1,2, Yue Yu1 Ning Ding1,2, Xingtai Lv1,2, Kai Tian1,2, Bowen Zhou1,∗ 1Department of Electronic Engineering, Tsinghua University, Beijing, China 2Frontis.AI, Beijing, China hem23@mails.tsinghua.edu.cn, zhoubowen@tsinghua.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable powerful potential across various downstream tasks after pre-training on large-scale corpora. However, their instruction-following skills and trustworthiness still fall short of expectations. Therefore, algorithms such as Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) (Ziegler et al., 2019; Ouyang et al., 2022; Lee et al., 2023) are utilized to further enhance LLMs’ abilities and align them better with human preferences.
@@
Considering the limited effectiveness of SFT and the high cost of data construction and training computation for RLHF, these two methods are often combined to leverage their respective strengths. Unfortunately, they can only be implemented as a sequential recipe constrained by the paradigm gap between SFT and early RLHF methods, stemming from differences in loss functions, data formats, and the requirement for auxiliary models.
@@
Figure 1: Comparison of Alignment Methods. IFT conducts alignment solely relying on positive samples and a single policy, starting from a pre-trained base model. IFT not only enjoys the second highest efficiency only lower than SFT, but also showcases comparable or even superior performance to other alignment algorithms.
Recently, Direct Preference Optimization (DPO) (Rafailov et al., 2024) was proposed to integrate Reward Modeling and Policy Optimization into one single procedure using a loss function derived from Proximal Policy Optimization (PPO) (Schulman et al., 2017), which demonstrates the potential to unify SFT and RLHF for the first time.
@@
Henceforth, many extended methods have been tried to realize this objective by bridging the gap between SFT and DPO. Some of them (Ethayarajh et al., 2024; Hong et al., 2024) aim to transform the contrastive loss of DPO into a SFT-like cross-entropy loss, learning positive samples similar to SFT while unlearning negative samples resort to Unlikelihood Training (Welleck et al., 2019). Some others get rid of the preference-labeling process before training, switching to collect samples and labels/rewards in an online manner (Liu et al., 2023; Yuan et al., 2024; Guo et al., 2024; Calandriello et al., 2024), or just treating the SFT targets and online policy generations as positive and negative samples respectively (Chen et al., 2024; Mitra et al., 2024; Liu et al., 2024). Nevertheless, preference-labeled pairwise data is still essential in these methods, and the need for reference model only becomes unnecessary in some cases. Thus the core differences between SFT and RLHF are not eliminated thoroughly. To address this challenging issue, a deeper and more unified understanding of SFT and RLHF are needed.
@@
In this paper, we attempt to explain the similarities and differences between SFT and RLHF by defining Preference Estimation and Transition Optimization in terms of state-action pairs within the Markov Decision Process (MDP) framework. Through this modeling, we demonstrate that SFT is simply a specialized case with inferior estimation and optimization among all RLHF methods. To be specific, classical RLHF methods usually sample the entire answer of policy based solely on the initial instruction, and align the preference reflected by the sampled sentence with human. In contrast, SFT only samples a single token based on intermediate state of target answer, which leads to a biased estimation of policy preference and an inferior alignment shortcut. Or put it in a metaphorical view, when preparing for an exam, it is obviously more beneficial to review reference answers after completing entire questions, rather than trying to infer each subsequent step based on the reference answer ahead and checking it immediately. It is the similar case in LMs’ training so that RLHF is more beneficial than SFT for enhancing LMs.
@@
Depending on this understanding, we introduce a unified alignment algorithm named Intuitive Fine-Tuning (IFT), which builds policy preference estimation in an intuitive manner like human, who can have a vague sense of the complete answer after hearing a question. Through a closer estimation to truly policy preference than SFT, IFT achieves a comparable or even superior alignment performance compared to the sequential recipe of SFT and RLHF. Additionally, IFT requires only a single policy model, and the same volume and format of data as SFT, enjoying both data and computation efficiency. These characteristics also enable IFT to be implemented in domains where preference data is unavailable or expensive to collect.
@@
Our main contribution are three folds: Firstly, by modeling within the MDP framework, we explain the similarities and differences between SFT and two fundamental RLHF methods, PPO and DPO. Secondly, building on this modeling, we introduce Intuitive Fine-tuning (IFT), a deeply unified version of SFT and RLHF, which enjoys similar efficiency to SFT, while having a closer performance to RLHF. Finally, through experiments in several benchmarks, we validate that IFT can perform comparably or even superiorly to different recipes of SFT and existing RLHF methods. In particular, IFT demonstrates significantly stronger reasoning and fact-following abilities than chosen baselines. An explainable toy-setting named Frozen Lake is also used for demonstrating the effectiveness of IFT.
@@ 2 Related Work
Classical Reinforcement learning (RL) has been demonstrated strong performance in various sequential decision-making and optimal control domains, including robotics (Levine et al., 2018), computer games (Vinyals et al., 2019) and others (Guan et al., 2021). There are two main categories of RL algorithms: value-based and policy-based, depending on whether they learn a parameterized policy. Value-based RL aims to fit an value function defined by Bellman Equation, containing methods such as Monte-Carlo (MC) Learning (Lazaric et al., 2007) and Temporal Difference Learning Sutton (1988); Seijen and Sutton (2014). However, value-based methods struggle in continuous or large discrete space for its greedy objective. Thus, policy-based methods were introduced to model the decision-making process using a parameterized policy. As one of its best-known algorithms, Proximal Policy Optimization (PPO) (Schulman et al., 2017) is widely used in various domains, including Natural Language Processing (NLP).
@@
Alignment for LMs has emerged as a crucial task these years, which adjusts the LMs’ generation distribution in line with human preferences. Building on the Bradley-Terry (BT) (Bradley and Terry, 1952) Model, current alignment methods typically train models to distinguish positive and negative answers given the same instruction using RLHF/RLAIF (Ziegler et al., 2019; Ouyang et al., 2022; Lee et al., 2023). While PPO remains the primary algorithm for alignment, its high demands for computation and memory hinders its broader use. As a result, DPO (Rafailov et al., 2024) proposes training a model to serve both as a policy model and a reward model, unifying reward modeling and policy training into a single process. Without sacrificing performance, DPO decrease the costly consumption of PPO through directly value iteration similar to a preference-based format of MC instead of TD. Nevertheless, the expensive preference-labeling process and the memory-hogging pair-wise data is still required.
@@
Improved Versions of DPO come out one after another. Efforts such as (Liu et al., 2023; Yuan et al., 2024; Guo et al., 2024; Calandriello et al., 2024; Chen et al., 2024; Mitra et al., 2024; Liu et al., 2024) focus on online sampling and automated label/reward collection, reducing the manual cost required for alignment. Methods like (Ethayarajh et al., 2024; Hong et al., 2024) aim to reduce DPO’s dependency on SFT warm-up by transforming its loss functions and data format into a SFT manner. These algorithms handle positive and negative samples using SFT objective and Unlikelihood Training (Welleck et al., 2019), respectively. However, the actual volume of training data is not decreased in these methods. Also, GPU-memory-consuming pair-wise data is still required, while the need for a reference model and preference-labeling for the entire answer trajectory is only eliminated in limited cases.
@@ 3.1 MDP in Language Models
The MDP applied to LMs can be formally described as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},r,\rho_{0})$, where $\mathcal{S}$ is the state space comprising ordered permutations of vocabularies, $\mathcal{A}$ is the action space consisting of vocabularies defined by the tokenizer, $\mathcal{T}$ is the transition matrix indicating token generation probabilities for a given state, $r$ represents rewards for state-action pairs, and $\rho_{0}$ is the initial state typically based on given instructions. See more details in Appendix A.1.
The primary objective of Language Modeling is to train a policy $\pi_{\theta}$ with $\mathcal{T}_{\theta}$ to mimic a human policy $\pi^{*}$ with $\mathcal{T}^{*}$, aiming for the two transition matrix to become identical:
@@
where $T$ is equivalence to $\mathcal{T}$, but instead, indicating the transition probability from one state to another.
@@ 3.2 Preference Estimation
$\mathcal{P_{\theta}}(\rho_{0})\rightarrow\mathcal{P^{*}}(\rho_{0})$ (4)
$\mathcal{P_{\theta}}(\rho_{0}):\rho_{0}\rightarrow[\pi_{\theta}(\rho_{0}),\pi_{\theta}(s_{1}^{\theta}),\pi_{\theta}(s_{2}^{\theta}),\ldots]$ (5)
$\mathcal{P^{*}}(\rho_{0}):\rho_{0}\rightarrow[\pi^{*}(\rho_{0}),\pi^{*}(s_{1}^{*}),\pi^{*}(s_{2}^{*}),\ldots]$ (6)
@@
To make preference optimizable, every policy’s preference can also be expressed as follows:
@@
Here, $\mathcal{S}_{\rho_{0}}$ denotes a conditional state space that constrained by the initial state $\rho_{0}$, within which each state can only be initially derived from $\rho_{0}$. Consequently, the model preference can be optimized through transition matrix, named Transition Optimization by us.
@@ 3.3 Transition Optimization
However, considering the limited data, only matrix elements representing state-action/state-state pairs contained in the dataset would be aligned. Given a target sample with instruction $\rho_{0}$ and length-$N$ answer, the objective would be:
$\forall a\in\mathcal{A},n\in[0,N]:\mathcal{T}_{\theta}(a,s_{n}^{*})\rightarrow\mathcal{T^{*}}(a,s_{n}^{*})$ (10)
@@
$\forall n\in[0,N]:T_{\theta}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$ (11)
$T_{\theta}(s_{n}^{*}|\rho_{0})=\prod\limits_{i=0}^{n}\mathcal{T}_{\theta}(a_{i}^{*}|s_{i}^{*})$ (12)
$T^{*}(s_{n}^{*}|\rho_{0})=\prod\limits_{i=0}^{n}\mathcal{T^{*}}(a_{i}^{*}|s_{i}^{*})$ (13)
where $s_{0}^{*}=\rho_{0}$ and $T^{*}(\rho_{0}|\rho_{0})=T_{\theta}(\rho_{0}|\rho_{0})=1$.
Consequently, the loss function can be derived from the disparities of the transition matrices between model and human.
3.4 From SFT to RLHF
We reformulate SFT, PPO and DPO using the aforementioned framework, detailed in table 1. A more comprehensible version is presented in figure 2. To compare the differences between them, we begin by introducing a fundamental theorem and corollary:
@@ 3.4 From SFT to RLHF
Table 1: Reformulation of SFT, PPO and DPO
Theorem Given a set of events $\mathcal{Z}$, the probability of any event $z\in\mathcal{Z}$ is between 0 and 1:
$\forall z\in\mathcal{Z}:0\leq P(z)\leq 1$ (14)
If all events are mutually independent, the sum of their probabilities equals 1:
$1=\sum_{z\in\mathcal{Z}}P(z)$ (15)
The event $z^{*}$ with the highest probability has a probability greater than or equal to any other event:
$\forall z\in\mathcal{Z}:0\leq P(z)\leq P(z^{*})\leq 1$ (16)

### v2

# Intuitive Fine-Tuning: Towards Simplifying Alignment into a Single Process

@@ Abstract
Supervised Fine-Tuning (SFT) and Preference Optimization (PO) are two fundamental processes for enhancing the capabilities of Language Models (LMs) post pre-training, aligning them better with human preferences. Although SFT advances in training efficiency, PO delivers better alignment, thus they are often combined. However, common practices simply apply them sequentially without integrating their optimization objectives, ignoring the opportunities to bridge their paradigm gap and take the strengths from both. To obtain a unified understanding, we interpret SFT and PO with two sub-processes — Preference Estimation and Transition Optimization — defined at token level within the Markov Decision Process (MDP) framework. This modeling shows that SFT is only a specialized case of PO with inferior estimation and optimization. PO evaluates the quality of model’s entire generated answer, whereas SFT only scores predicted tokens based on preceding tokens from target answers. Therefore, SFT overestimates the ability of model, leading to inferior optimization. Building on this view, we introduce Intuitive Fine-Tuning (IFT) to integrate SFT and Preference Optimization into a single process. IFT captures LMs’ intuitive sense of the entire answers through a temporal residual connection, but it solely relies on a single policy and the same volume of non-preference-labeled data as SFT. Our experiments show that IFT performs comparably or even superiorly to sequential recipes of SFT and some typical Preference Optimization methods across several tasks, particularly those requires generation, reasoning, and fact-following abilities. An explainable Frozen Lake game further validates the effectiveness of IFT for getting competitive policy. Our code will be released at https://github.com/TsinghuaC3I/Intuitive-Fine-Tuning.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable powerful potential across various downstream tasks after pre-training on large-scale corpora brown2020language ; achiam2023gpt ; zhou2024generative . However, their instruction-following skills and trustworthiness still fall short of expectations bender2021dangers ; bommasani2021opportunities ; li2022trust . Therefore, algorithms such as Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) ziegler2019fine ; ouyang2022training ; lee2023rlaif are used to further enhance LLMs’ abilities and align them better with human preferences.
@@
Considering the limited effectiveness of SFT and the high cost of data construction and training computation for RLHF, these two methods are often combined to leverage their respective strengths. Unfortunately, they are typically implemented as a sequential recipe constrained by the paradigm gap between SFT and early RLHF methods, stemming from differences in loss functions, data formats, and the requirement for auxiliary models.
@@
Figure 1: Comparison of Alignment Methods. IFT conducts alignment solely relying on positive samples and a single policy, starting from a pre-trained base model. IFT not only shows high efficiency as SFT, but also performs comparably or even superiorly to other alignment algorithms.
@@
Recently, Direct Preference Optimization (DPO) rafailov2024direct was proposed to integrate Reward Modeling and Policy Optimization into one single procedure using a loss function derived from Proximal Policy Optimization (PPO) schulman2017proximal . This approach demonstrates the potential to unify SFT and RLHF for the first time. Henceforth, many extended methods have been tried to realize this objective by bridging the gap between SFT and DPO. Some of them ethayarajh2024kto ; hong2024reference ; zhang2024negative aim to transform the contrastive loss of DPO into a SFT-like cross-entropy loss, learning positive samples similar to SFT while unlearning negative samples resort to Unlikelihood Training welleck2019neural . Some others get rid of the preference-labeling process before training, switching to collect samples and labels/rewards in an online manner liu2023statistical ; yuan2024self ; guo2024direct ; calandriello2024human ; tajwar2024preference , or just treating the SFT targets and online policy generations as positive and negative samples respectively xiong2023iterative ; chen2024self ; mitra2024orca ; liu2024extensive . Nevertheless, preference-labeled pairwise data is still essential in these methods, and the need for reference model only becomes unnecessary in some cases. Thus the core differences between SFT and Preference Optimization are not eliminated thoroughly. To address this challenging issue, a deeper and more unified understanding of them are needed.
@@
In this paper, we attempt to explain the similarities and differences between SFT and Preference Optimization by defining Preference Estimation and Transition Optimization in terms of state-action pairs within the Markov Decision Process (MDP) framework. Through this modeling, we demonstrate that SFT is simply a specialized case with inferior estimation and optimization among all Preference Optimization methods. To be specific, typical Preference Optimization methods usually sample the entire answer of policy based solely on the initial instruction, and align the policy preference reflected by the sampled sentence with human preferences. In contrast, SFT only samples a single token based on intermediate state of target answer, which leads to a biased estimation of policy preference and an inferior alignment shortcut. Or put it in a metaphorical view, when preparing for an exam, it is obviously more beneficial to review reference answers after completing entire questions, rather than trying to infer each subsequent step based on the reference answer ahead and checking it immediately. It is the similar case in LMs’ training, where SFT is less beneficial for enhancing LMs.
@@
Depending on this understanding, we introduce a unified alignment algorithm named Intuitive Fine-Tuning (IFT), which builds policy preference estimation in an intuitive manner like human, who can have a vague sense of the complete answer after hearing a question. Through a closer estimation to truly policy preference than SFT, IFT achieves a comparable or even superior alignment performance compared to the sequential recipe of SFT and Preference Optimization. Additionally, IFT requires only a single policy model, and the same volume and format of data as SFT, enjoying both data and computation efficiency. These characteristics also enable IFT to be implemented in domains where preference data is unavailable or expensive to collect.
@@
Our main contribution are three folds: 1) By modeling within the MDP framework, we explain the similarities and differences between SFT and two fundamental Preference Optimization methods, PPO and DPO; 2) Based on this modeling, we introduce Intuitive Fine-tuning (IFT), a deeply unified version of SFT and Preference Optimization. IFT utilizes temporary residual connections to extract the model’s generation preference given the initial instructions, offering similar efficiency to SFT while achieving performance closer to Preference Optimization; 3) Through experiments on several benchmarks, we validate that IFT performs comparably or even superiorly to various SFT recipes and existing Preference Optimization methods. In particular, IFT demonstrates significantly stronger reasoning and fact-following abilities than the chosen baselines. An explainable toy-setting Frozen Lake further demonstrates the effectiveness of IFT.
@@ 2 Related Work
Classical Reinforcement learning (RL) has demonstrated strong performance in various sequential decision-making and optimal control domains, including robotics levine2018learning , computer games vinyals2019grandmaster and others guan2021direct . There are two main categories of RL algorithms: value-based and policy-based, depending on whether they learn a parameterized policy. Value-based RL aims to fit an value function defined by Bellman Equation, containing methods such as Monte-Carlo (MC) Learning lazaric2007reinforcement and Temporal Difference Learning sutton1988learning ; seijen2014true . However, value-based methods struggle in continuous or large discrete space for its greedy objective. Thus, policy-based methods were introduced to model the decision-making process using a parameterized policy. As one of its best-known algorithms, Proximal Policy Optimization (PPO) schulman2017proximal is widely used in various domains, including Natural Language Processing (NLP).
@@
Alignment for LMs has emerged as a crucial task these years, which adjusts the LMs’ generation distribution in line with human preferences. Building on the Bradley-Terry (BT) bradley1952rank Model, current alignment methods typically train models to distinguish positive and negative answers given the same instruction using RLHF/RLAIF ziegler2019fine ; ouyang2022training ; lee2023rlaif . While PPO remains the primary algorithm for alignment, its high demands for computation and memory hinders its broader use. Consequently, many improved methods have been proposed dong2023raft ; yuan2023rrhf ; zhao2023slic . Among them, DPO rafailov2024direct unifies reward modeling and policy optimization by utilizing a loss function derived from PPO, training a single model to serve as both a policy model and a reward model. Without sacrificing performance, DPO decrease the costly consumption of PPO through directly value iteration similar to a preference-based format of MC instead of TD. Nevertheless, the expensive preference-labeling process is still required. Additionally, a warm-up stage based on SFT is also necessary, which may bring trade-offs in fitting different objectives between SFT and Preference Optimization.
@@
Improved Versions of DPO come out one after another. Efforts such as liu2023makes ; khaki2024rs ; yin2024relative ; guo2024controllable ; bansal2024comparing ; liu2024extensive try to enhance the contrastive learning by utilizing better ranking strategies, more informative data, or more number of negative samples. except for using offline data, liu2023statistical ; yuan2024self ; guo2024direct ; calandriello2024human ; chen2024self ; mitra2024orca focus on online sampling and automated label/reward collection, reducing the manual cost required for alignment. Recently, methods like ethayarajh2024kto ; hong2024reference aim to reduce DPO’s dependency on SFT warm-up by transforming its loss functions and data format into a SFT manner. These algorithms handle positive and negative samples using SFT objective and Unlikelihood Training welleck2019neural , respectively. However, the actual volume of training data is not decreased in these methods. Also, GPU-memory-consuming pair-wise data is still required, while the need for a reference model and preference-labeling for the entire answer trajectory is only eliminated in limited cases.
@@ 3.1 MDP in Language Models
The MDP applied to LMs can be formally described as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},r,\rho_{0})$, where $\mathcal{S}$ is the state space comprising ordered permutations of vocabularies, $\mathcal{A}$ is the action space consisting of vocabularies defined by the tokenizer, $\mathcal{T}$ is the transition matrix indicating token generation probabilities for given states, $r$ represents rewards for state-action pairs, and $\rho_{0}$ is the initial state typically based on given instructions. See more details in Appendix A.1.
The primary objective of Language Modeling is to train a policy $\pi_{\theta}$ with $\mathcal{T}_{\theta}$ to mimic a human policy $\pi^{*}$ with $\mathcal{T}^{*}$, aiming for the two transition matrices to become identical:
@@
where $T$ is equivalence to $\mathcal{T}$, but instead, indicating the transition probability between states.
@@ 3.2 Preference Estimation
$\mathcal{P_{\theta}}(\rho_{0})\rightarrow\mathcal{P^{*}}(\rho_{0}):\bigg{\{}\begin{aligned} &\mathcal{P_{\theta}}(\rho_{0}):\rho_{0}\rightarrow[\pi_{\theta}(\rho_{0}),\pi_{\theta}(s_{1}^{\theta}),\pi_{\theta}(s_{2}^{\theta}),\ldots]\\ &\mathcal{P^{*}}(\rho_{0}):\rho_{0}\rightarrow[\pi^{*}(\rho_{0}),\pi^{*}(s_{1}^{*}),\pi^{*}(s_{2}^{*}),\ldots]\end{aligned}$ (4)
@@
To make preference optimizable, the policy’s preference can also be expressed as follows:
@@
Here, $\mathcal{S}_{\rho_{0}}$ denotes a conditional state space that constrained by the initial state $\rho_{0}$, within which each state can only be initially derived from $\rho_{0}$. Consequently, the model preference can be optimized through transition matrix, named Transition Optimization.
@@ 3.3 Transition Optimization
However, considering the limited data, only matrix elements representing state-action/state-state pairs contained in the dataset $\mathcal{D}$ would be aligned. Given a data sample with instruction $\rho_{0}$ and target answer with length-$N$, the objective would be:
$\forall a\in\mathcal{A},n\in[0,N],\rho_{0}\in\mathcal{D},s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*}:\mathcal{T}_{\theta}(a,s_{n}^{*})\rightarrow\mathcal{T^{*}}(a,s_{n}^{*})$ (8)
@@
$\forall n\in[0,N],\rho_{0}\in\mathcal{D},s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*}:T_{\theta}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$ (9)
$T_{\theta}(s_{n}^{*}|\rho_{0})=\prod\limits_{i=0}^{n}\mathcal{T}_{\theta}(a_{i}^{*}|s_{i}^{*}),~{}T^{*}(s_{n}^{*}|\rho_{0})=\prod\limits_{i=0}^{n}\mathcal{T^{*}}(a_{i}^{*}|s_{i}^{*})$ (10)
where $s_{0}^{*}=\rho_{0}$, $T^{*}(\rho_{0}|\rho_{0})=T_{\theta}(\rho_{0}|\rho_{0})=1$, and $s_{i}^{*}$ denotes the intermediate state of target answer.
Consequently, the loss function can be derived from the disparities of the transition matrices between model and human. Some typical loss function are listed in Appendix A.4
Table 1: Reformulation of SFT, PPO and DPO
@@
3.4 From SFT to Preference Optimization
We reformulate SFT, PPO and DPO using the aforementioned framework, detailed in Table 1 and Appendix A.4. A more comprehensible version is presented in Figure 2. To compare the differences between them, we begin by introducing a fundamental theorem and corollary:
Theorem Given a set of events $\mathcal{Z}$, the probability of any event $z\in\mathcal{Z}$ is between 0 and 1, i.e., $\forall z\in\mathcal{Z}:0\leq P(z)\leq 1$. If all events are mutually independent, the sum of their probabilities equals 1, i.e., $1=\sum_{z\in\mathcal{Z}}P(z)$. The event $z^{*}$ with the highest probability has a probability greater than or equal to any other event, i.e., $\forall z\in\mathcal{Z}:0\leq P(z)\leq P(z^{*})\leq 1$.